OpenStudy (anonymous):

Solve the triangle. A = 19°, C = 102°, c = 6

OpenStudy (anonymous):

i got B = 59°, a ≈ 2, b ≈ 5.3

OpenStudy (anonymous):

just checking my answer

OpenStudy (anonymous):

59 is of course correct

OpenStudy (anonymous):

\[\frac{6}{\sin(102)}=\frac{a}{\sin(19)}\] so \[a=\frac{6\sin(19)}{\sin(102)}\] is that about 2?

OpenStudy (anonymous):

yeah looks like about 2

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

want me to check the last one? probably right as well

OpenStudy (anonymous):

no thts okay thank u

OpenStudy (wolf1728):

I previously entered some figures here but I deleted those because I was thinking (incorrectly) this was two angles and an included side. 59° side 6 19° IS an ASA congruency problem and the remaining sides are: 5.2579 1.997